Consider an annuity of payments of 2.500$ at the end of every second year. What is the present value of this annuity if it runs for ten years and the interest rate is j1 = 7%?


1
Expert's answer
2021-11-01T07:17:25-0400

"PV=C\cdot \frac{1-(1+i)^{-n}}{i}"

where C is cash flow per period,

i is interest rate,

n is number of payments


"PV=2500\cdot \frac{1-(1+0.07)^{-5}}{0.07}=10250.49\$"